OpenStudy (nottim):

IF anyone is still awake: sin(7pi/12)+cos(pi/12)

OpenStudy (nottim):

I should know this, But erk.

OpenStudy (anonymous):

7pi/12*180/pi pi/12*180/pi cos 15 + sin 105

OpenStudy (nottim):

Oops. Sorry. I did not clarify: Simplify the following. the final answer must be exact. However, it isn't necessary to rationalize the denominator.

OpenStudy (anonymous):

use angle identites for cos (45-30) and sin (60+45)

OpenStudy (nottim):

Sorry. I don't understand....

OpenStudy (anonymous):

sincos sincos cos cos sin sin??

OpenStudy (nottim):

Oh.

OpenStudy (anonymous):

NotHello NotTim, sin(7pi/12) + cos(pi/12) cos(pi/2 + pi/12) + cos (pi/2)

OpenStudy (anonymous):

NotSorry, sin(pi/2 + pi/12) + cos(pi/2)

OpenStudy (nottim):

Er, I still don't understand...

OpenStudy (nottim):

Wait. So am I doing the addition formula for both of them?

OpenStudy (nottim):

In that, I break both of them up?

OpenStudy (anonymous):

Don;t do it my way, it's only gonna make it more complicated I thought I would be able to reduce it to Zero but Zero isn't the answer... shinigam1m is right!

OpenStudy (anonymous):

sin(105) + cos(15) sin(60 + 45) + cos(45 - 30) sin60cos45 + sin45cos60 + cos45cos30 + sin45sin30

OpenStudy (nottim):

I think I m suppose to keep it in radians.

OpenStudy (anonymous):

keep it in radians then 45 = pi/4, 60 = pi/3 and 30=pi/6

OpenStudy (anonymous):

i changed my picture